OpenStudy (anonymous):

Is it possible to solve d((x+1)/(x-1))/dx? I reached here halfway through a question

OpenStudy (anonymous):

Ah!! That's right! I completely forgot about it!

OpenStudy (anonymous):

What, ans is zero?

OpenStudy (anonymous):

@Dio no, I just couldn't differentiate it before. I tried splitting it into two fractions. but i remember how to do it now. =D

OpenStudy (anonymous):

Use the quotient rule \[\frac {1(x-1) - 1(x+1)} {(x-1)^2}= \frac {-2} {(x-1)^2} \]